# Channel Measurement, Modeling, and Simulation for 6G: A Survey and Tutorial

### v3

@@ III-B3 Channel sparsity
When THz waves impinge upon “rough” surfaces, they give rise to diffuse scattering phenomena. In comparison to low-frequency specular reflections, diffuse scattering-generated multipath energy is weaker. This leads to a considerable reduction in the quantity and power of NLOS MPCs received at the receiver, resulting in most power concentrated on a few MPCs, namely, sparsity.
@@
The channel sparsity can be expressed in both the angular and delay domains and in the low rank of the channel matrix, which refers to a small number of coefficients containing a large proportion of the energy [146]. It is often claimed that mmWave channels are “sparse,” i.e., have few entries in the delay angle bins (ITU-R M.2412 [42]). The sparse mmWave channel at 28 GHz has been discovered and studied in 2018 [147]. So, due to the weak diffraction ability, the THz channel is more likely to be sparse due to the weak diffraction ability [16, 148]. The THz channel is more likely to be sparse due to the weak diffraction ability [16, 148]. [149] quantified the sparsity of the channel using the Gini index. Channel measurements conducted in sub-6 GHz, mmWave, and THz bands confirmed the sparsity characteristic in the THz channel. Then, an intracluster power allocation model within clusters was proposed to model the sparsity in the delay domain. [72] analyzed the number of clusters of delay, angle and power domains. The cluster numbers were reduced, which reflected the sparsity of the THz channel.
III-C THz Channel Modeling
@@
Existing THz channel modeling approaches can be broadly classified as stochastic, deterministic, and hybrid approaches [150]. THz stochastic channel models are mainly obtained based on measurements and describe the type of environment rather than a specific location. While parameterizing such models from measurements requires significant effort, implementing channel modeling from stochastic models is much less complex than implementing deterministic models [151].
@@
THz deterministic approaches solve (approximate) Maxwell’s system of equations in a given environment and can achieve high accuracy [152], but they require detailed information about the geometric and electromagnetic properties of the environment and have high computational complexity, such as RT method [153]. RT-based THz channel modeling has been extensively utilized in sub-6 GHz and mmWave bands [154]. However, channel modeling using the RT approach for the THz band is different from that in the sub-6 GHz and mmWave bands. These differences result in new opportunities for RT-based channel modeling in THz bands. For the inherent characteristics of RT, RT based on geometric optics can be explained by the high-frequency approximation of Maxwell’s equations. Consequently, the quasioptical characteristics make the RT results in the THz band more reliable and suitable than those in the sub-6 GHz band. Additionally, for THz channel characteristics, the high propagation loss and sparsity characteristics of the THz channel enhance the need for site-specific analysis, which can be well captured by RT simulations.
@@
However, there are still some challenges for RT-based channel modeling in THz bands. Most notably, there is no complete electromagnetic property of the material in THz bands [155]. Additionally, there is a lack of adequate measurements for validating and calibrating RT [148]. For real-time applications, the RT channel model is difficult to implement. Therefore, a data and model dual-driven channel modeling method is presented, which uses channel measurement data to calibrate simulation parameters [26]. The data and model dual-driven channel simulation flow is shown in Fig. 8. The material EM properties in RT simulations are tuned to reach the best agreement in terms of power and delay for the dominant propagation paths. In the beginning, the RT simulation outputs the results corresponding to the initial EM parameters, which are compared with the measurement data. If the results of the power and delay parameters for the dominant paths match (with the objective to minimize the RMS error), the simulation results are output. If the results do not match, the relative permittivity and conductivity of the materials are updated, and the simulation results are output again. This operation is repeated until the best agreement between the simulation and measurement results is achieved.
@@
The performance of the data and model dual-driven THz channel modeling method is demonstrated through a comparison between simulations and channel measurements. Fig. 9 shows the measurement power angle profile in which the five dominant paths are marked by numbers. Furthermore, the delay of the five dominant paths is plotted to perform a comparison. Fig. 9 shows the results of RT simulations for the five dominant paths. By comparing Fig. 9 with Fig. 9, it can be seen that MPCs have good agreement in both the spatial and delay domains between the measurement and RT results. For example, the differences between the measurements and RT simulations for the delay of the five paths are within $\pm$2 ns. The maximum differences in the angle of arrival, angle of departure, delay, and power gain of the dominant paths are $-1.8^{\circ}$, $-1.5^{\circ}$, 0.7 ns and -0.7 dB, respectively. Therefore, the data and model dual-driven channel modeling method can describe well the propagation characteristics, i.e., the delay and spatial dispersion, with reduced simulation complexity in the THz bands [26].
Figure 9: Data and model dual-driven channel modeling method results: a) Power, angle of arrival, and angle of departure measurement results; b) Power, angle of arrival, and angle of departure simulation results [26].
@@
THz hybrid channel models combine deterministic and stochastic channel modeling approaches, providing a balance between accuracy and efficiency, such as in map-based channel models [156] and hybrid ray and graph models [157]. For example, due to the sparsity of the THz channel, the main propagation paths of the THz channel (e.g., LOS path and primary reflection path) are generated by the deterministic model, while the other multipaths are generated by the stochastic model, which ensures the accuracy of the main propagation paths and quickly obtains the remaining multipath information [158]. Therefore, the hybrid channel model is a compromised approach for THz channel modeling.
III-D Summary and Prospects
@@
In summary, extensive THz channel measurement campaigns have been carried out in typical scenarios, such as InH, data centers, and outdoor streets. Existing studies indicate that the THz band exhibits significantly different channel characteristics compared to sub-6 GHz and mmWave bands, including characteristics such as reflection, high path loss, and sparsity. Stochastic, deterministic, and hybrid channel modeling approaches have been employed in THz channel modeling. Among them, the data and model dual-driven THz channel modeling method show promising prospects. Despite the research on THz channels in recent years, many open issues motivate future research efforts in the following areas [25].
III-D1 THz channel measurement platforms with higher performance
@@
Current THz channel measurement platforms are unable to achieve a balance between measurement accuracy, measurement speed, and measurement distance, which constrains the study of THz channel characteristics. In addition, existing virtual MIMO THz channel measurements based on rotating horn antennas are time-consuming. Therefore, it is necessary to build a THz E-MIMO array and phased array to achieve accurate parameter estimation and fast beam sweeping.
III-D2 THz channel characteristics

### v4

@@ V-C4 Channel sparsity
When THz waves impinge upon “rough” surfaces, they give rise to diffuse scattering phenomena. Compared with low-frequency specular reflections, diffuse scattering-generated multipath energy is weaker. This leads to a considerable reduction in the quantity and power of NLOS MPCs received at the receiver, resulting in most of the power being concentrated on a few MPCs, also known as sparsity.
@@
The channel sparsity can be expressed in both the angular and delay domains and in the low rank of the channel matrix, which refers to a small number of coefficients containing a large proportion of the energy [229]. It is often claimed that mmWave channels are “sparse,” i.e., have few entries in the delay angle bins (ITU-R M.2412 [46]). The sparse mmWave channel at 28 GHz was discovered and studied in 2018 [230]. Therefore, due to its weak diffraction ability, the THz channel is more likely to be sparse [16, 231]. [51] quantified the sparsity of a channel via the Gini index. Channel measurements conducted in sub-6 GHz, mmWave, and THz bands confirmed the sparsity characteristic in the THz channel. Then, an intracluster power allocation model within clusters was proposed to model the sparsity in the delay domain. [137] analyzed the number of clusters of delay, angle and power domains. The number of clusters was reduced, which reflected the sparsity of the THz channel.
V-D THz Channel Modeling
@@
Existing THz channel modeling approaches can be broadly classified into stochastic, deterministic, and hybrid approaches [232]. THz stochastic channel models are obtained mainly on the basis of measurements and describe the type of environment rather than a specific location. While parameterizing such models from measurements requires significant effort, implementing channel modeling from stochastic models is much less complex than implementing deterministic models [233].
@@
THz deterministic approaches solve (approximatly) Maxwell’s system of equations in a given environment and can achieve high accuracy [234], but they require detailed information about the geometric and electromagnetic properties of the environment and have high computational complexity, such as the RT method [235]. RT-based THz channel modeling has been extensively utilized in sub-6 GHz and mmWave bands [236]. However, channel modeling using the RT approach for the THz band is different from that for the sub-6 GHz and mmWave bands. These differences result in new opportunities for RT-based channel modeling in THz bands. For the inherent characteristics of RT, RT based on geometric optics can be explained by the high-frequency approximation of Maxwell’s equations. Consequently, the quasioptical characteristics make the RT results in the THz band more reliable and suitable than those in the sub-6 GHz band. Additionally, for THz channel characteristics, the high propagation loss and sparsity characteristics of the THz channel increase the need for site-specific analysis, which can be well captured by RT simulations.
@@
However, there are still some challenges for RT-based channel modeling in THz bands. Most notably, there is no complete electromagnetic property of the material in THz bands [237]. Additionally, there is a lack of adequate measurements for validating and calibrating RT [231]. For real-time applications, the RT channel model is difficult to implement. Therefore, a data and model dual-driven channel modeling method is presented, which uses channel measurement data to calibrate simulation parameters [31]. The data and model dual-driven channel simulation workflow is shown in Fig. 20. The material EM properties in RT simulations are tuned to reach the best agreement in terms of power and delay for the dominant propagation paths. Initially, the RT simulation outputs the results corresponding to the initial EM parameters, which are compared with the measurement data. If the results of the power and delay parameters for the dominant paths match (with the objective of minimizing the RMS error), the simulation results are output. If the results do not match, the relative permittivity and conductivity of the materials are updated, and the simulation results are output again. This operation is repeated until the best agreement between the simulation and measurement results is achieved.
@@
The performance of the data and model dual-driven THz channel modeling method is demonstrated through a comparison between simulations and channel measurements. Fig. LABEL:THz_RT_1 shows the measurement power angle profile in which the five dominant paths are marked by numbers. Furthermore, the delays of the five dominant paths are plotted for comparison. Fig. LABEL:THz_RT_2 shows the results of RT simulations for the five dominant paths. By comparing Fig. LABEL:THz_RT_1 with Fig. LABEL:THz_RT_2, it can be seen that the MPCs have good agreement in both the spatial and delay domains between the measurement and RT results. For example, the differences between the measurements and RT simulations for the delay of the five paths are within $\pm$2 ns. The maximum differences in the angle of arrival, angle of departure, delay, and power gain of the dominant paths are $1.8^{\circ}$, $1.5^{\circ}$, 0.7 ns and -0.7 dB, respectively. Therefore, the data and model dual-driven channel modeling method can effectively describe the propagation characteristics, i.e., the delay and spatial dispersion, with reduced simulation complexity in the THz bands [31].
@@
THz hybrid channel models combine deterministic and stochastic channel modeling approaches, providing a balance between accuracy and efficiency, such as in map-based channel models [238] and hybrid ray and graph models [239]. For example, due to the sparsity of the THz channel, the main propagation paths of the THz channel (e.g., the LOS path and primary reflection path) are generated by the deterministic model, whereas the other multipaths are generated by the stochastic model, which ensures the accuracy of the main propagation paths and quickly obtains the remaining multipath information [240]. Therefore, the hybrid channel model is a compromise approach for THz channel modeling.
V-E Summary and Prospects
New mid-band channel measurement campaigns have been carried out in typical scenarios, such as UMa, UMi, and InH. The channel characteristics of the new mid-band are different from those of other frequency bands such as sub-6 GHz, mmWave and THz bands. The path loss is less than that in high-frequency bands such as mmWave and THz, which is conducive to coverage. The delay spread is less than sub-6 GHz, and a smaller protection interval can be set in the frame structure. The new mid-band combines the advantages of coverage and capacity, making it the major candidate frequency band for future 6G applications.
@@ V-E Summary and Prospects
Existing studies indicate that the THz band exhibits different channel characteristics based on THz band channel measurements and characteristic comparisons, including characteristics such as scattering, reflection, high path loss, and sparsity. Stochastic, deterministic, and hybrid channel modeling approaches have been employed in THz channel modeling. Among them, the data and model dual-driven THz channel modeling methods show promising prospects. Despite the research on THz channels in recent years, many open issues motivate future research efforts in the following areas [30].
V-E1 Channel measurement platforms with higher performance
@@
Current channel measurement platforms are unable to achieve a balance between measurement accuracy, measurement speed, and measurement distance, which constrains the study of THz channel characteristics. In addition, existing virtual MIMO channel measurements based on rotating horn antennas are time-consuming. Therefore, it is necessary to build a multi-frequency XL-MIMO array and phased array to achieve accurate parameter estimation and fast beam sweeping.
V-E2 Channel characteristics